News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 25 November 2005 7.00am EST

Happy Thanksgiving from the Chandra Team!

During the last week Chandra completed the observing schedule as planned.

A realtime procedure was completed on November 18 to update the on-board Ephemeris.

A realtime procedure was completed on November 22 to uplink a flight software patch to increase the value of the Electron Proton Helium INstrument (EPHIN) E1300 threshold by a factor of two. The new value will reduce the number of false triggers of the science instrument safing sequence (SCS 107) during periods of puffed-up radiation belts, and when EPHIN is at high temperature. The change will also allow for the relaxation of a mission planning thermal constraint.

A second flight software patch was uplinked on November 22 to modify SCS107 to execute only a single move of the Science Instrument Module (SIM). The previous implementation consisted of two moves, the first to an intermediate position to allow the HRC camera door to be closed, and a second to its final HRC-S position. The change was made since the HRC is no longer closed during SCS 107 runs, and the two moves can now cause excessive SIM motor temperature levels.

The schedule of targets for next week is shown below and includes a series of observations of GRS1915+105 coordinated with RXTE.

------------------------------------------
SDSS J073739.96+38      ACIS-S       Nov 28
IC 0396                 ACIS-S       Nov 29
SDSS J1552+3201         ACIS-S
Abell 1650              ACIS-I
UGC 08041               ACIS-S
A478_OFFSET2            ACIS-I
HD93129AB               ACIS-S/HETG
Radiation Belts
EGS-6                   ACIS-I       Nov 30
GRS1915+105             ACIS-S/HETG
A478                    ACIS-I       Dec 1
HD93129AB               ACIS-S/HETG
GRS1915+105             ACIS-S/HETG
A478 (2 obs)            ACIS-I
MWC 147                 ACIS-S
HD93129AB               ACIS-S/HETG  Dec 2
SDSS J0209-0005         ACIS-S
Radiation Belts
EGS-6 (2 obs)           ACIS-I       Dec 3
GRS1915+105             ACIS-S/HETG
A478                    ACIS-I
NGC 0949                ACIS-S
NGC 0959                ACIS-S
EGS-6                   ACIS-I       Dec 4
PSR B1929+10            ACIS-S
A478                    ACIS-I
PKS 0920-397            ACIS-S
EGS-6                   ACIS-I
SDSS J0735+2659         ACIS-S       Dec 5

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports